# Constructing a T-test for Value Function Comparison of Individualized Treatment Regimes in the Presence of Multiple Imputation for Missing Data

### v1

@@ 4.1 CHNS: Data Analysis
Hypertension, also known as high blood pressure, is a chronic medical condition that affects over one billion adults worldwide (Bloch, 2016). It is a major risk factor for cardiovascular disease, which is the leading cause of death globally (Bromfield and Muntner, 2013). People with hypertension are more likely to develop kidney disease (Weldegiorgis and Woodward, 2020), vision problems (Bhargava et al., 2012), and cognitive impairment (Kilander et al., 1998). Hypertension can also affect quality of life and increase annual medical costs (Wang et al., 2017). Over the past two to three decades, China has experienced a significant increase in the prevalence of hypertension (from 20.8% in 2004 to 29.6% in 2010 and 24.7% in 2018) due to increased life expectancy and lifestyle changes (Zhang et al., 2023; Wang et al., 2023). As hypertension has become a significant burden on China’s population health, it is important to study this condition and develop effective prevention and treatment strategies. While medication proves effective in managing hypertension, the general adherence to anti-hypertensive medication remains low. This can be attributed to factors such as a low reimbursement ratio, high daily medical costs, and access to healthcare Cui et al. (2020). In addition to taking medications, adopting non-pharmacologic interventions such as engaging in regular physical activity, reducing salt intake, and following a balanced diet rich in fruits and vegetables can help to reduce high blood pressure (Appel, 2003). However, there is potential heterogeneity in which individuals may see the most benefit from behavioral interventions.
@@
Our overall objective is to identify the most effective non-pharmacologic treatments for subpopulations of individuals, considering the challenges of adhering to a comprehensive healthy lifestyle based on public health recommendations. As a first step, we examine the potential variability in the effectiveness of exceeding 5 hours of weekly moderate-to-vigorous physical activity (MVPA) in reducing the risk of hypertension in the CHNS population. We seek to distinguish between participants who benefit from engaging in MVPA for more than 5 hours per week and those who do not. For participants who see no hypertension benefit from increasing their physical activity, it is possible to identify other potential interventions to consider, such as diet modification. Understanding which interventions confer optimal outcomes empowers individuals to identify the most effective and efficient behavioral treatments and their associated benefits. This knowledge can be highly beneficial, as individuals are more motivated to engage in behaviors when they are aware of their positive outcomes Cane et al. (2012). To address this research question, we tested whether a personalized physical intervention strategy would improve population-level health outcomes compared to a one-size-fits-all intervention.
@@
We used data from the China Health and Nutrition Survey (CHNS)Zhang et al. (2014), a population-based, observational data set consisting of high-quality data on diet (3-day repeated 24-hour recalls) and physical activity (detailed 7-day recall) collected from individuals in China using detailed recall instruments. Our analysis focused on the study year 2009 with 8320 adults. Our outcome, hypertension, was defined  [^1] Anti-hypertensive medication is included in the model in our analysis, so blood pressure measurements here are not adjusted for medication. Participants who were taking medication may exhibit varying responses to changes in physical activity. Consequently, we have included medication as a covariate that has the potential to influence how individuals respond to engaging in more than 5 hours of physical activity per week. as a systolic blood pressure $\geq$ 130 mmHg or diastolic blood pressure $\geq$ 80 mmHg. Individuals with missing systolic or diastolic blood pressure measurements or who were missing more than 10% of their covariates were excluded from our study, resulting in a final analytic sample of 5241 individuals, 60% of whom had hypertension. In the CHNS dataset, physical activity encompassing leisure, occupational, transportation, and domestic activities, were collected via a comprehensive survey recall instrument Ng et al. (2014). Our study focused on weekly moderate-to-vigorous physical activity (MVPA), defined as activities with a minimum of 3 METS (Metabolic equivalent of task) Ainsworth et al. (2011)  [^2] The METS, or Metabolic equivalent of task, is a unit of measurement for physical activity. One MET is equivalent to a person’s oxygen consumption at a rate of 3.5 milliliters per kilogram per minute. Sylvia et al. (2014). Our sample population has a high MVPA level, with a median of 15 hours per week (Ng et al., 2014). Targeting the 5-hour MVPA weekly guideline (Olson et al., 2019; Piercy et al., 2018; Ng et al., 2014), we dichotomized the treatment into a binary variable $A$: over 5 hours ($A=1$) and 5 or fewer hours ($A=0$) of MVPA per week. In the CHNS data, 66% of the sample population engaged in more than 5 hours of MVPA weekly. With insights from biological understanding and published literature Ng et al. (2014), we identified twenty risk factors for hypertension. These factors also have the potential to influence the relationship between physical activity and hypertension, including: age, gender, caloric intake, BMI, education, smoking status, sodium intake, potassium intake, alcohol consumption, hypertension medication, household income, province, and urbanization index. Less than 15% of our sample had missing covariates, handled using missForest package in R for multiple imputation (Stekhoven, 2022; Stekhoven and Buehlmann, 2012). Table 2 provides further details about these factors based on one imputed dataset. The other imputed datasets exhibit similar distributions.
@@
Figure 3: The value function results of six different treatment regimes: observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR. For comparison purposes, we include the average outcome as a benchmark. (a) shows the results on original data, and (b)(c)(d) shows the three simulated data with three levels of treatment effect modification $\delta=1$, $\delta=2$, $\delta=10$, respectively.
@@
Figure 4: The value function differences among selected pairs of treatment regimes. Specifically, the Q-learning optimal ITR comparison with the observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, and D-learning optimal ITR are shown. (a) shows the results on original data, and (b)(c)(d) shows the three simulated data with three levels of treatment effect modification $\delta=1$, $\delta=2$, $\delta=10$, respectively.
@@
The data was split into a 70% training set and a 30% test set, with missForest imputation performed 10 times on each set to avoid dependency. We obtained Q-learning optimal ITRs (Qian and Murphy, 2011) and D-learnings optimal ITR (Tian et al., 2014) for each of the training sets using a logistic regression model with elastic-net penalization. We then compared the value function of each treatment rule on the test set, including the observed treatment, assigning all individuals to treatment $A=1$ or $A=0$, optimal ITRs derived from Q-learning and D-learning. The value functions of these treatment rules were estimated using the common estimator $V$ and the differences between value functions were directly obtained by subtracting the value functions and then averaged over 10 imputed data sets. The variance of the value function and the variance of the difference between the value functions were estimated by our method, which accounts for both modeling and multiple imputation variance.
Table 3: Estimated value functions of different treatment regimes for CHNS data.
Training Set Test Set
Mean SD Mean SD
Obs 0.602 0.021 $0.595$ 0.031
AllLowPA 0.603 0.018 $0.591$ 0.026
AllHighPA 0.600 0.011 $0.598$ 0.017
Q-Learning 0.568 0.013 $0.574$ 0.017
D-Learning 0.576 0.016 $0.598$ 0.023
Table 4: Estimated value differences and p-values of selected comparisons between treatment regimes for CHNS data.
Training Set Test Set
Difference Mean SD P-Value Mean SD P-Value
AllLowPA vs. Obs 0.001 0.011 0.924 -0.004 0.017 0.827
AllHighPA vs. Obs -0.001 0.018 0.953 0.003 0.026 0.897
AllHighPA vs. AllLowPA -0.002 0.021 0.921 0.007 0.032 0.821
Q vs. Obs -0.033 0.017 0.046 -0.021 0.026 0.429
Q vs. AllLowPA -0.034 0.017 0.040 -0.017 0.026 0.519
Q vs. AllHighPA -0.032 0.013 0.014 -0.024 0.018 0.176
Q vs. D -0.008 0.013 0.562 -0.023 0.020 0.239
@@
Our goal is to compare personalized and uniform interventions in reducing hypertension risk by comparing the value functions in test sets for (1) observed physical activity, under the assumption that individuals continue their existing routines; (2) assigning all individuals to treatment $A=1$, assuming all individuals are above the 5 hours weekly MVPA recommendation; (3) assigning all individuals to treatment $A=0$, assuming all individuals are at or below the 5 hours weekly MVPA recommendation; and (4) optimal ITRs derived from various methods. Figure 3(a) shows the value function and variances for those treatment regimes. The Q-learning value function is the lowest. This means that if the population follows the Q-learning ITR, the expected risk of hypertension in the population is lower than that if the population follows the other treatment assignment, although the differences between the value functions are small. Figure 4(a) shows the pairwise difference of the value functions between treatment regimes that we are interested in. Our focus is on determining the necessity of recommending over 5 hours of weekly MVPA to everyone or specific subgroups. Thus we compare the best ITR, Q-learning ITR, defined by the lowest value function estimate in the training set (Table 3) with the one-size-fits-all treatment $A=1$ for the test set. Based on the t-test result, the value functions are not significantly different, with a p-value of 0.18 for the Q-learning personalized treatment rule compared to assigning all individuals to treatment $A=1$ (Table 4). This suggests that a personalized approach does not significantly outperform the population-level physical activity recommendations for doing physical activity for more than five hours per week. Although the training data results show that the Q-learning has a significantly lower value function than the best one-size-fits-all treatment $A=1$, the results shown here are test set results. The test set results hold particular significance as they underscore the model’s generalizability. They offer insight into the performance expected of the trained model when confronted with new and unseen data.
@@
Our findings suggest that adjustments in physical activity for a subset of individuals alone do not yield a substantial reduction in hypertension risk at the population level. Alternative non-pharmacologic interventions may demonstrate greater efficacy in mitigating hypertension risk across specific subpopulations. This variance in response may be influenced by the interaction between multiple factors, such as sodium intake and physical activity. For instance, individuals with high sodium consumption may not benefit significantly from increased physical activity alone. In contrast, sodium intake reduction could potentially offer a more effective strategy for lowering hypertension risk. Another possible explanation is that much physical activity in the CHNS data comprised working, as opposed to leisure, activities. Published data suggest that occupational physical activity may not be as protective against hypertension Huai et al. (2013). The complete pairwise comparison of all the comparisons is depicted in the appendix Figure 6.
4.2 CHNS: Augmented Analysis
These differences across all the value functions in our original CHNS analysis could truly be this small in magnitude. However, it is important to note that these models only consider a small subset of factors that could impact how changes in physical activity can impact hypertension. Hypertension is complex, resulting from an interplay between individual behavior, environment, and additional factors like metabolism and genetics. These factors might help provide additional important insight and better identify underlying heterogeneity in the impact of physical activity on hypertension not captured in our original analysis. Therefore, we wanted to investigate our method’s performance in the presence of prescriptive treatment effect interaction, which corresponds to these additional factors, or combinations of these additional factors with our existing factors, that explain more of the variation in how an individual’s hypertension risk would be impacted by changes in physical activity. Therefore, an additional analysis was carried out on simulated datasets with modified prescriptive treatment effect interaction terms.
The outcome variables in these data sets were generated by the fitted Q-learning outcome regression model with the treatment effect interaction terms (excluding the treatment main effect term) multiplied by a factor of $\delta\geq 1$. We investigate three options of $\delta$ for this analysis: $\delta=1$, $\delta=2$, and $\delta=10$, representing the estimated original, double, and tenfold treatment effect, respectively. In each analysis, hypertension outcomes are replaced by generated outcomes from the modified Q-learning outcome regression model. This augmented analysis examines our method’s performance when the outcome generative process is known and how the performance varies as the true prescriptive treatment effect enlarges.
@@ 4.2 CHNS: Augmented Analysis
As $\delta$ increases, from $\delta=1$, to $\delta=2$, to $\delta=10$ (Figure 3 (b)-(d)), the value function differences between the Q-learning strategy and other treatment rules increase. Figure 4 shows the difference in magnitude between the pairwise differences in value function results between the Q-learning treatment rules and both (1) assigning all individuals to either treatment $A=1$ and (2) treatment $A=0$ across our original results. The complete pairwise comparison of all the comparisons is depicted in the appendix figure 6. In all these instances, the Q-learning models deliver the best ITRs, with values notably lower, suggesting a lower risk of hypertension in the treatment-effect-augmented population especially when compared to population interventions when all individuals are assigned to low physical activity or all individuals are assigned to high physical activity. When $\delta=1$, we observe that the pattern of value functions bears resemblance to that of the original data across various methods. As the prescriptive treatment effect becomes increasingly evident, both Q-learning and D-learning demonstrate progressively superior performance compared to the ’one-size-fits-all’ approaches. This underscores the importance of personalizing over population-level interventions when certain factors significantly affect individual responses to treatment. Applying these methods to existing rich, detailed observational data can provide an important insight into factors that might distinguish these subgroups of responders. Future work will build on these existing models to better incorporate additional high-dimensional data. However, these results suggest that if a large heterogeneous effect truly exists, our method can successfully capture it.
@@ 5 Discussion
In this paper, we propose an innovative t-test based approach that can directly compare the value functions of any two treatment regimes. The validity of the approach follows from the asymptotic normality of the standard value function for the estimated treatment regime and the asymptotic normality of the propensity score model parameters. Our method provides valid estimates for (1) the variance for a value function for a single treatment regime, (2) the variance for the difference between two value functions of two treatment regimes, (3) the p-value of the t-test for the significance of the difference between two value functions, and (4) the application of these estimations in scenarios involving multiple imputations. This method maintains simplicity in variance calculation and is computationally more efficient than the bootstrap method, especially when multiple imputation is used for handling missing data. Through simulation studies and the data application example, we demonstrate the performance and the ease of implementation of our method in different scenarios. Additionally, this method enables the evaluation and comparison of ITR effectiveness using abundant observational cohort data when clinical trial data are not available. This comparison is crucial as individuals seek behavioral modification guidance amidst numerous recommendations, and emphasizing targeting strategies could increase the likelihood of effective changes.
Nevertheless, it is important to acknowledge a limitation inherent in our methodology – its reliance on the presupposition that the propensity score model holds true. In observational data, extreme propensity scores can influence the estimates of the value function as well as the variance estimates of the value function. For future work, it will be interesting to use more robust estimators to alleviate the impact of potential misspecification of the propensity score model. Another extension of this method is to multiple-stage ITR comparison. Expanding our method to simultaneously compare more than two ITRs would also present an interesting avenue for further exploration.

### v3

@@ 4.1 CHNS: Data Analysis
Hypertension, also known as high blood pressure, is a chronic medical condition that affects over one billion adults worldwide 19. It is a major risk factor for cardiovascular disease, which is the leading cause of death globally 20. People with hypertension are more likely to develop kidney disease, 21 vision problems, 22 and cognitive impairment. 23 Hypertension can also affect quality of life and increase annual medical costs 24. Over the past two to three decades, China has experienced a significant increase in the prevalence of hypertension (from 20.8% in 2004 to 29.6% in 2010 and 24.7% in 2018) due to increased life expectancy and lifestyle changes 25, 26. As hypertension has become a significant burden on China’s population health, it is important to study this condition and develop effective prevention and treatment strategies. While medication proves effective in managing hypertension, the general adherence to anti-hypertensive medication remains low. This can be attributed to factors such as a low reimbursement ratio, high daily medical costs, and access to healthcare 27. In addition to taking medications, adopting non-pharmacologic interventions such as engaging in regular physical activity, reducing salt intake, and following a balanced diet rich in fruits and vegetables can help to reduce high blood pressure 28. However, there is potential heterogeneity in which individuals may see the most benefit from behavioral interventions.
@@
Our overall objective is to identify the most effective non-pharmacologic treatments for subpopulations of individuals, considering the challenges of adhering to a comprehensive healthy lifestyle based on public health recommendations. As a first step, we examine the potential variability in the effectiveness of exceeding 5 hours of weekly moderate-to-vigorous physical activity (MVPA) in reducing the risk of hypertension in the CHNS population. We seek to distinguish between participants who benefit from engaging in MVPA for more than 5 hours per week and those who do not. For participants who see no hypertension benefit from increasing their physical activity, it is possible to identify other potential interventions to consider, such as diet modification. Understanding which interventions confer optimal outcomes empowers individuals to identify the most effective and efficient behavioral treatments and their associated benefits. This knowledge can be highly beneficial, as individuals are more motivated to engage in behaviors when they are aware of their positive outcomes 29. To address this research question, we tested whether a personalized physical activity intervention strategy could achieve population-level health outcomes comparable to an intervention that assigns high physical activity to everyone.
@@
We used data from the China Health and Nutrition Survey (CHNS)30, a population-based, observational data set consisting of high-quality data on diet (3-day repeated 24-hour recalls) and physical activity (detailed 7-day recall) collected from individuals in China using detailed recall instruments. Our analysis focused on the study year 2009 with 8320 adults. Our outcome, hypertension, was defined  [^4] Anti-hypertensive medication is included in the model in our analysis, so blood pressure measurements here are not adjusted for medication. Participants who were taking medication may exhibit varying responses to changes in physical activity. Consequently, we have included medication as a covariate that has the potential to influence how individuals respond to engaging in more than 5 hours of physical activity per week. as a systolic blood pressure $\geq$ 130 mmHg or diastolic blood pressure $\geq$ 80 mmHg. Individuals with missing systolic or diastolic blood pressure measurements or who were missing more than 10% of their covariates were excluded from our study, resulting in a final analytic sample of 5241 individuals, 60% of whom had hypertension. In the CHNS dataset, physical activity encompassing leisure, occupational, transportation, and domestic activities, were collected via a comprehensive survey recall instrument 31. Our study focused on weekly moderate-to-vigorous physical activity (MVPA), defined as activities with a minimum of 3 METS (Metabolic equivalent of task) 32  [^5] The METS, or Metabolic equivalent of task, is a unit of measurement for physical activity. One MET is equivalent to a person’s oxygen consumption at a rate of 3.5 milliliters per kilogram per minute. 33. Our sample population has a high MVPA level, with a median of 15 hours per week 31. Targeting the 5-hour MVPA weekly guideline 2, 1, 31, we dichotomized the treatment into a binary variable $A$: over 5 hours ($A=1$) and 5 or fewer hours ($A=0$) of MVPA per week. In the CHNS data, 66% of the sample population engaged in more than 5 hours of MVPA weekly. With insights from biological understanding and published literature 31, we identified twenty risk factors for hypertension. These factors also have the potential to influence the relationship between physical activity and hypertension, including: age, gender, caloric intake, BMI, education, smoking status, sodium intake, potassium intake, alcohol consumption, hypertension medication, household income, province, and urbanization index. Less than 15% of our sample had missing covariates, handled using missForest package in R for multiple imputation 34, 35. Table 4 provides further details about these factors based on one imputed dataset. The other imputed datasets exhibit similar distributions. The data was split into a 70% training set and a 30% test set, with missForest imputation performed 10 times on each set to avoid dependency. We obtained Q-learning optimal ITRs 4 and D-learnings optimal ITR 5 for each of the training sets using a logistic regression model with elastic-net penalization. We then compared the value function of each treatment rule on the test set, including the observed treatment, assigning all individuals to treatment $A=1$ or $A=0$, optimal ITRs derived from Q-learning and D-learning.
@@
Figure 3: The value function and 95% confidence interval estimates of six different treatment regimes: observed treatment (Obs), lower physical activity hours (AllLowPA, $A=0$) for all individuals, higher physical activity hours (AllHighPA, $A=1$) for all individuals, Q-learning optimal ITR (Q), and D-learning optimal ITR (D). For comparison purposes, we include the average outcome as a benchmark. (a) shows the results on original data, and (b)(c)(d) shows the three simulated data with three levels of treatment effect modification $\delta=1$, $\delta=2$, $\delta=10$, respectively.
@@
Figure 4: The value function differences and 95% confidence interval among selected pairs of treatment regimes. Specifically, the Q-learning optimal ITR comparison with the observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, and D-learning optimal ITR are shown. (a) shows the results on original data, and (b)(c)(d) shows the three simulated data with three levels of treatment effect modification $\delta=1$, $\delta=2$, $\delta=10$, respectively.
@@
The value functions of these treatment rules were estimated using the common estimator $V$ and the differences between value functions were directly obtained by subtracting the value functions and then averaged over 10 imputed data sets. The variance of the value function and the variance of the difference between the value functions were estimated by our method, which accounts for both modeling and multiple imputation variance.
@@
Our goal is to compare personalized and uniform interventions in reducing hypertension risk by comparing the value functions in test sets for (1) observed physical activity, under the assumption that individuals continue their existing routines; (2) assigning all individuals to treatment $A=1$, assuming all individuals are above the 5 hours weekly MVPA recommendation; (3) assigning all individuals to treatment $A=0$, assuming all individuals are at or below the 5 hours weekly MVPA recommendation; and (4) optimal ITRs derived from various methods. Figure 3(a) shows the value function and 95% confidence interval for those treatment regimes. The differences between the value functions for different decision rules are small. Especially, the Q-learning value function is the lowest but the confidence interval intersects with the that of other treatment regimes. This means that if the population follows the Q-learning ITR, the expected risk of hypertension in the population is similar to when the population follows the other treatment assignment (e.g. high PA for everyone). Figure 4(a) shows the pairwise difference of the value functions between treatment regimes that we are interested in. Our focus is on determining the necessity of recommending over 5 hours of weekly MVPA to everyone or specific subgroups. Thus we compare the best ITR, Q-learning ITR, defined by the lowest value function estimate in the training set (Table S1) with the one-size-fits-all treatment $A=1$ for the test set. Based on the t-test result, the value functions are not significantly different, with a p-value of 0.156 for the Q-learning personalized treatment rule compared to assigning all individuals to treatment $A=1$ (Table S2). This suggests that a personalized approach can achieve a similar hypertension risk as the population-level physical activity recommendations for doing physical activity for more than five hours per week.The results presented herein pertain to the test set, which is essential for evaluating the model’s generalizability to new, unseen data. A comparison of model performance between the training and test sets is provided in Appendix Figure S4 and Figure S5. The training set results closely mirror those of the test set, underscoring the model’s stability; however, the test set results are emphasized as they more accurately reflect the model’s anticipated performance in real-world applications.
@@
Our findings indicate that applying a uniform high level of physical activity across the entire population does not significantly reduce hypertension risk compared to tailoring activity levels to individual characteristics. Specifically, our results suggest that increased physical activity may benefit certain subgroups, whereas others do not derive additional benefits beyond 5 hours of weekly MVPA. For these subpopulations, alternative non-pharmacologic interventions may be more effective in mitigating hypertension risk. For instance, individuals with high sodium consumption may not benefit significantly from increased physical activity alone. In contrast, sodium intake reduction could potentially offer a more effective strategy for lowering hypertension risk. Another possible explanation is that much physical activity in the CHNS data comprised working, as opposed to leisure, activities. Published data suggest that occupational physical activity may not be as protective against hypertension 36. The complete pairwise comparison of all the comparisons is depicted in the appendix Figure S5.
4.2 CHNS: Augmented Simulation Analysis
In Section 4.1, our method did not detect a significant difference between the two treatment rules under comparison. This finding prompted the question of whether the null result was attributable to limitations of the method when applied to real-world data, or if the true difference was indeed minimal. To address this issue, we conducted an augmented analysis to evaluate our method’s performance in scenarios where a true difference between the treatment rules is present. It is important to note that the models in Section 4.1 considered only a limited set of factors influencing the effect of physical activity on hypertension. Given the complexity of hypertension—arising from the interplay of individual behavior, environmental influences, and additional factors such as metabolism and genetics—incorporating these variables could reveal important insights and better identify the underlying heterogeneity in the impact of physical activity on hypertension. Consequently, we examined our method’s performance in the presence of prescriptive treatment effect interactions that account for these additional factors, or their combinations with our existing factors, which may explain more of the variation in individual hypertension risk. An additional analysis was therefore conducted using simulated datasets with modified prescriptive treatment effect interaction terms.
@@ 4.2 CHNS: Augmented Simulation Analysis
The outcome variables in these data sets were generated by the fitted Q-learning outcome regression model with the treatment effect interaction terms (excluding the treatment main effect term) multiplied by a factor of $\delta\geq 1$. We investigate three options of $\delta$ for this analysis: $\delta=1$, $\delta=2$, and $\delta=10$, representing the estimated original, double, and tenfold treatment effect, respectively. In each analysis, hypertension outcomes are replaced by generated outcomes from the modified Q-learning outcome regression model. This augmented analysis examines our method’s performance when the outcome generative process is known and how the performance varies as the true prescriptive treatment effect enlarges. As $\delta$ increases, from $\delta=1$, to $\delta=2$, to $\delta=10$ (Figure 3 (b)-(d)), the value function differences between the Q-learning strategy and other treatment rules increase. Figure 4 shows the difference in magnitude between selected pairwise differences in value function results between the Q-learning treatment rules, assigning all individuals to either treatment $A=1$, and treatment $A=0$.  [^6] The complete pairwise comparison of all the comparisons is depicted in the appendix Figure S5. Across all instances, the Q-learning models produced the most effective ITRs, with values decreasing significantly as $\delta_{t}$ increased. This trend indicates a reduced risk of hypertension in the heterogeneous, treatment-effect-augmented population, especially compared to population-level interventions where all individuals are assigned to high physical activity. When $\delta=1$, we observe that the pattern of value functions bears resemblance to that of the original data across various methods. As the prescriptive treatment effect becomes increasingly evident, both Q-learning and D-learning demonstrate progressively superior performance compared to the ’one-size-fits-all’ approaches. This underscores the importance of personalizing over population-level interventions when certain factors significantly affect individual responses to treatment. Applying these methods to existing rich, detailed observational data can provide an important insight into factors that might distinguish these subgroups of responders. Future work will build on these existing models to better incorporate additional high-dimensional data. However, these results suggest that if a large heterogeneous effect truly exists, our method can successfully capture it.
@@ 5 Discussion
In this paper, we propose an innovative t-test based approach that can directly compare the value functions of any two treatment regimes. The validity of the approach follows from the asymptotic normality of the standard value function for the estimated treatment regime and the asymptotic normality of the propensity score model parameters. Our method provides valid estimates for (1) the variance and 95% confidence interval for a value function for a single treatment regime, (2) the variance and 95% confidence interval for the difference between two value functions of two treatment regimes, (3) the p-value of the t-test for the significance of the difference between two value functions, and (4) the application of these estimations in scenarios involving multiple imputations. Our method’s estimates exhibit comparable Type I error and power performance to those obtained via bootstrap. Moreover, our approach simplifies variance calculation and is computationally more efficient, especially when multiple imputation is used to handle missing data. Through simulation studies and the data application example, we demonstrate the performance and the ease of implementation of our method in different scenarios. Additionally, this method enables the evaluation and comparison of ITR effectiveness using abundant observational cohort data when clinical trial data are not available. This comparison is crucial as individuals seek behavioral modification guidance amidst numerous recommendations, and emphasizing targeting strategies could increase the likelihood of effective changes.
Nevertheless, it is important to acknowledge a limitation inherent in our methodology – its reliance on the presupposition that the propensity score model holds true. In observational data, extreme propensity scores can influence the estimates of the value function as well as the variance estimates of the value function. For future work, it will be interesting to use more robust estimators to alleviate the impact of potential misspecification of the propensity score model. Another extension of this method is to multiple-stage ITR comparison. Extending our method to simultaneously compare more than two ITRs presents an interesting avenue for future research, which would require establishing the asymptotic normality of the joint propensity scores for multiple treatment assignments.